OpenStudy (gabylovesyou):

Describe how to transform the quantity of the third root of x to the fourth power, to the fifth power into an expression with a rational exponent. Make sure you respond with complete sentences.

OpenStudy (gabylovesyou):

\[\left( \sqrt[3]{x ^{4}} \right) ^{5}\]

OpenStudy (anonymous):

hi gaby what grade are u in

OpenStudy (gabylovesyou):

10th :) @mathsciencehistory

OpenStudy (anonymous):

lol im in 11th and i dont know that are you in honors or AP classes

OpenStudy (gabylovesyou):

im advanced.. so it should be 11th grade math..

OpenStudy (gabylovesyou):

@hartnn

OpenStudy (anonymous):

.-. moderators are not good for me

OpenStudy (anonymous):

@iGreen or @Jaynator495

OpenStudy (anonymous):

one of them can help

OpenStudy (gabylovesyou):

thanks lol ^-^

hartnn (hartnn):

\(\huge \sqrt[n]{x^m} = x^{(\dfrac{m}{n})}\)

hartnn (hartnn):